OpenStudy (anonymous):

I need to find an equation of the line having the slope and the given point. I cant seem to get the right answer and ive been working on it forever: m=4/5, (1,-6)

OpenStudy (anonymous):

(y+6) = 4/5 *(x-1) => 5y +30 = 4x -4 => 4x -5y = 34

OpenStudy (anonymous):

Im still confused I understand the above part which is y+6= 4/5 (x-1) but I need the final answer to be y = and then the equation

OpenStudy (anonymous):

y = 4/5 x - 34/5

OpenStudy (anonymous):

both are same......... different form..............

OpenStudy (anonymous):

okay thanks. I get so confused when fractions are involved, ive never been good at it

OpenStudy (anonymous):

if u realy want help to me ............ you are welcome to talk to me using skype........ sudhanshu_kmr

OpenStudy (anonymous):

okay well i think thats all i needed though, thanks!